OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-1, 5) m = -1 i got y=4?

OpenStudy (anonymous):

b= 4 not y

OpenStudy (anonymous):

y = mx + b 5 = 1 + b b = 4 y = -x + 4

OpenStudy (anonymous):

y = -x + 4